OpenStudy (anonymous):

They gave me two points and asked me to find out an equation of the line that passes through the points. The points are (5,-1) , (-5,5).

OpenStudy (anonymous):

1. Find the slope. 2. Find the y intercept.

OpenStudy (anonymous):

1. Slope equals \[ \frac{-1-5}{5-(-5)} = -\frac{3}{5} \]

OpenStudy (anonymous):

2. y intercept equals 2

OpenStudy (anonymous):

So the equation of the line is y=-3/5 x + 2

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(5--1)/(-5-5) m=(6)/(-5-5) m=(6)/(-10) m=-3/5 So the slope of the line that goes through the points is m=-3/5 Now let's use the point slope formula: y-y1=m(x-x1) y--1=(-3/5)(x-5) y+1=(-3/5)(x-5) y+1=(-3/5)x+(-3/5)(-5) y+1=(-3/5)x+3 y=(-3/5)x+3-1 y=(-3/5)x+2 So the equation is \[\Large y=-\frac{3}{5}x+2\]

OpenStudy (anonymous):

Thank you!